A mixture of polarised & unpolarised is incident on a polariser. As the polariser is rotated through 360°, it is found that the minimum & maximum of the transmitted intensity is in the ratio 1:q. The ratio of intensities of polarised to unpolarised light in the incident beam is =?
1
Expert's answer
2019-08-08T10:56:41-0400

In the case of an unpolarized light beam, 50% of the flux always passes through the polarizer

"I_1=I_{10}\cdot 0.5"

In the case of a polarized light beam, its intensity varies according to the Malus law:

"I_2=I_{20}\cos^2(\phi)"

Then the maximum and minimum of the past intensity:

"I_{max}=\frac{I_{10}}{2}+I_{20}\\
I_{min}=\frac{I_{10}}{2}"

Knowing their relationship, we find:

"\frac{1}{q}=\frac{\frac{I_{10}}{2}}{\frac{I_{10}}{2}+I_{20}}=\frac{I_{10}}{I_{10}+2I_{20}}\\
I_{10}+2I_{20}=qI_{10}\\
2I_{20}=(q-1)I_{10}\\
\frac{I_{10}}{I_{20}}=\frac{2}{1-q}"